OpenStudy (anonymous):

medels can someone please help 3/3^-2

jhonyy9 (jhonyy9):

hope you know that a^-2 = 1/(a^2) yes ?

OpenStudy (anonymous):

no clue

jhonyy9 (jhonyy9):

but there is a^x ---- = a^(x-y) a^y can using this ?

OpenStudy (anonymous):

i don't understand you

jhonyy9 (jhonyy9):

so this is equal 3^(1-(-2)) = 3^(1+2) = 3^3

jhonyy9 (jhonyy9):

do you know the property of exponents ?

OpenStudy (anonymous):

yeah but thats not 1 of my choices

jhonyy9 (jhonyy9):

why ? 3 not equal 3^1 ?

OpenStudy (anonymous):

i thought it was 1/3 but idk

jhonyy9 (jhonyy9):

so there is 3^1 ------ = 3^(1-(-2)) = 3^(1+2) 3^(-2) ok ?

OpenStudy (anonymous):

i no i got 3/1 but the anser isn't in the choices so i thought it was 1/3

jhonyy9 (jhonyy9):

what are the choices ?

OpenStudy (anonymous):

they are 27 18 1/2 or 1/3

OpenStudy (anonymous):

4 choices

jhonyy9 (jhonyy9):

27 is the right answer because 3^3 =27 ok ?

OpenStudy (anonymous):

ok thx

jhonyy9 (jhonyy9):

do you understand it now sure ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

because 3x3x3

jhonyy9 (jhonyy9):

ok was my pleasure

jhonyy9 (jhonyy9):

yes right

OpenStudy (anonymous):

yup i made 100%